OpenStudy (kainui):

What is the integral:

OpenStudy (kainui):

\[\Large \int\limits \frac{dx}{\sqrt{1-x^2}}\] Show your work. =)

hartnn (hartnn):

i would have started it by trying \(\large x = \sin u\)

hartnn (hartnn):

dx = cos u du \(\Large \int \dfrac{\cos u du }{\cos u}\)

hartnn (hartnn):

ugh, can't get any easier :P

hartnn (hartnn):

\(\int 1 du = u +c= \sin^{-1}x +c\)

OpenStudy (kainui):

Wait, here's where my issue comes in: \[\Large \int\limits \frac{\cos \theta}{|\cos \theta|} d \theta = \int\limits SquareWave(\theta) d \theta\]

OpenStudy (kainui):

 Maybe there's really no problem hmmm. Oh well lol.

OpenStudy (kainui):

Yeah never mind false alarm I guess I thought for some reason this would get us different answers but we can't unless we let x be complex.